Consider the reaction for the decomposition of hydrogen disulfide:

2H2S(g)⇌2H2(g)+S2(g),  
Kc = 1.67×10-7 at 800∘C

The reaction is carried out at the same temperature with the following initial concentrations:

[H2S]=3.20×10-4M

[H2]=0.00M

[S2]=0.00M

Find the equilibrium concentration of S2.

Express the concentration to three significant figures and include the appropriate units.
Answer #1

Equilibrium reaction is

2 H2S (g)  \rightarrow H2 (g) + S2 (g)

Now ICE table for the reaction is

[H2S] [H2] [S2]
initial 3.20*10-4 0.00 0.00
change -2x +2x +x
equilibrium (3.20*10-4 - 2x) 2x x

Now, Equilibrium constant (Kc ) = \frac{[H_{2}]^{2}[S]}{[H_{2}S]^{2}}

now, putting the value of concentration of H2S, H2 and S2 at equilibrium

Kc = \frac{(2x)^{2}\times x}{(3.20\times 10^{-4} -2x)^{2}}

Solving Cubic equation with this small data is very tough.

As, Equilibrium constant is very less, the reaction will not proceed very far towards right then x will be also very small

Therefore, (3.20*10-4 - 2x) \approx 3.20*10-4

now,

\frac{(2x)^{2}\times x}{(3.20\times 10^{-4})^{2}} = Kc

or , \frac{(2x)^{2}\times x}{(3.20\times 10^{-4})^{2}} = 1.67*10-7

or, 4x3 =  1.67*10-7 * (3.20*10-4)2

or, 4x3 = 17.10*10-15

or, x3 = (17.10/4) * 10-15

or, x3 = 4.27*10-15

or, x = 1.62*10-5 M

therefore

at equilibrium, [H2] = 2*x = 2*1.62*10-5 = 3.24*10-5 M

[S2] = x = 1.62*10-5 M

[H2S] = 3.20*10-4 - (2*1.62*10-5 ) M = 2.88*10-4 M

now, Kc = \frac{[H_{2}]^{2}[S]}{[H_{2}S]^{2}} =  \frac{(3.24*10^{-5})^{2}*1.62*10^{-5}}{(2.88*10^{-4})^{2}}

= 2 *10-7  

The calculated Kc value is nearly close to the given value of Kc , hence answer was valid.